OpenStudy (anonymous):

The acceleration of an object is given by the equation a(t)=6t^3-3t^2+2 m/s^2 Determine its velocity and position after 1 seconds, assuming s(0)=4m and v(0)= 0 m/s After 1 seconds, the object's velocity is______m/s and its position is___________m

OpenStudy (anonymous):

find the anti derivative of \[6t^3-3t^2+2\]

Parth (parthkohli):

Velocity is the anti-derivative of acceleration. Position is the anti-derivative of velocity.

OpenStudy (anonymous):

your answer will have a C in it, which you can solve for by setting \(v(0)=0\) and seeing that C must be zero as well

OpenStudy (anonymous):

3x^4/2-x^3+2x+c

OpenStudy (anonymous):

i can walk you through the steps if you like, it is all using \[\int t^n=\frac{t^{n+1}}{n+1}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3t^4/2-t^3+2t+c

OpenStudy (anonymous):

\[\frac{3}{2}t^4-t^3+2t+ c\]

OpenStudy (anonymous):

and since \(v(0)=0=c\) you also have \(c=0\) making it \[v(t)=\frac{3}{2}t^4-t^3+2t\]

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

then the first question asks for \(v(1)\) which is now real easy, replace \(t\) by \(1\)

OpenStudy (anonymous):

then find the anti derivative a second time to get \(s(t)\)

OpenStudy (anonymous):

so 5/2 will be v(1)

OpenStudy (anonymous):

and then 101/20

OpenStudy (anonymous):

i am not sure what you mean by "then 101/20"

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

can you help me with another problem?

OpenStudy (anonymous):

\(v(1)=\frac{5}{2}\) looks good to me

OpenStudy (anonymous):

but i did not find \(s(t)\) did you?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i got that for the second part i found the second dreviitve and i s(1) and c=4, it came up as the right answer

OpenStudy (anonymous):

although i am not sure you finished with this one did you find \(s(t)=\frac{3}{10}t^5-\frac{t^4}{4}+t^2+4\)?

OpenStudy (anonymous):

The marginal revenue for x items in dollars is given by r'(x)=-2.5x+7 determine and the revenue function and demand function

OpenStudy (anonymous):

oh i see you are right

OpenStudy (anonymous):

can you help me with the problem i posted please

OpenStudy (anonymous):

i am not sure about this one a quick google search tells me marginal revenue is \[R(x)=P(x)x\] does that look familiar?

OpenStudy (anonymous):

yes but im not sure if it has to do with this

OpenStudy (anonymous):

we have \[R'(x)=-2.5x+7\] so \[R(x)=-1.25x^2+7x\]

OpenStudy (anonymous):

if we write \[-1.25x^2+7=(-1.25x+7)x\] it looks like \[P(x)x\]

OpenStudy (anonymous):

i am not exactly sure either, sorry it looks like you want two things, revenue function and demand function maybe i can find it in a book

OpenStudy (anonymous):

thanks though im going to repost it to see if anyone can help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it looks like "revenue function" is \(R\) so in this case it is \[R(x)=-1.25x^2+7x\] and demand function is \(p(x)=-1.25x+7\) since \[R(x)=xP(x)\]